You purchased a bond at a price of $2,000. In 15 years when the bond matures, the bond will be worth $10,000. It is exactly 9 years after you purchased the bond and you can sell the bond today for $6,100. If you hold the bond until it matures, what annual rate of return will you earn from today?
Answer #1

r = [FV/PV]1/n - 1

= [$10,000/$2,000]1/15 - 1 = 1.1133 - 1 = 0.1133, or 11.33%